OpenStudy (anonymous):

MEDAL FOR HELP PLZ Which choice could be the equation of a line perpendicular to the line represented by this equation? Y=5x-2 A) y=-1/5x+5 B) y=5x+2 C) y=1/5x-7 D) y=-5x+5

OpenStudy (anonymous):

Perpendicular slopes: \[m_1 = -\frac{ 1 }{ m_2 }\]

OpenStudy (anonymous):

In a linear equation the coefficient of x represents the slope.

OpenStudy (anonymous):

I dont know what any of that means

OpenStudy (anonymous):

Do you know what a slope is?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

If two lines are perpendicular to each other, their slopes will be opposite and negative to each other

OpenStudy (anonymous):

x = 1/x opposite x = -x negative

OpenStudy (anonymous):

in your question, 5x is the slope

OpenStudy (anonymous):

what is the opposite and negative of 5x?